OpenStudy (anonymous):

It is known that the refractive index u of the material of a prism,depends on the wavelength as follows: u=a+b/λ^2 Where a,and b are constants.Plot a graph between u and λ. Main question : Identify the pair of variables to be used to get a straight line graph

OpenStudy (anonymous):

When plotted against lambda, it will curve, getting large at small lambda. Plotting y against (1/lambda^2) or plotting log y vs log lambda could produce linear (straight line) plots.

OpenStudy (anonymous):

how do i know which pair gives straight line? @douglaswinslowcooper

OpenStudy (anonymous):

Of course you could plot them. I chose them because they are of the form z = m x + b, a linear equation for z in variable x in the first suggestion z is y and x is (1/lambda^2) in the second, z is log y and x is log(1/lambda^2) = -2 log (lambda)

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

\(\large \mathbb{u=a+b/\lambda^2} \) \(\large \mathbb{u-a= b/\lambda^2} \) take log both sides

ganeshie8 (ganeshie8):

\(\large \mathbb{\log (u-a)= \log(b/\lambda^2)}\) \( \large \mathbb{\log (u-a)= \log b - \log \lambda^2}\) \( \large \mathbb{\log (u-a)= \log b - 2\log \lambda} \) put \(y = \log(u-a)\) \(x = \log \lambda \)

ganeshie8 (ganeshie8):

\(y = \log b - 2x\) straight line, wid slope = -2, y-intercept = log b

ganeshie8 (ganeshie8):

so the new change of variables wud be : log(u-a) and log(lambda)

OpenStudy (anonymous):

book mein Plotting u against (1/lambda^2) likha hai :/ ill prefer to know how did we get that combination !

ganeshie8 (ganeshie8):

that also works right ?

ganeshie8 (ganeshie8):

simply put 1/lambda^2 = x and stare at the equation

ganeshie8 (ganeshie8):

u=a+b/λ^2 u = a + bx straight line

ganeshie8 (ganeshie8):

oh you're asking, how to get to that analytically ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

let me think a bit, logifying is the dumb method i remember...

OpenStudy (anonymous):

okay forget the analytical method :P us wale se kaam chal jaega 8|

ganeshie8 (ganeshie8):

wait a sec, logifying is good when u have an exponent funtion like : u = b/2^λ

ganeshie8 (ganeshie8):

but wat you're having is simply a hyperbola... so just parameterizing x = 1/λ^2 will do i hope... sorry cant provide anything useful :|

OpenStudy (anonymous):

even if I have something like \[\Huge \mu=a+\frac{b}{\lambda^{100000}}\] i can put 1/lambda^100000000000 = x and it will still be a straight line??

ganeshie8 (ganeshie8):

ofcourse, if \(\mu\) and \(\lambda\) are the variables and \(a, b\) are constant, given equation becomes : \(\mu = a + bx\)

ganeshie8 (ganeshie8):

which is a linear equation in two variables : \(\mu , x\) produces a straight line

OpenStudy (anonymous):

that is cool...:o

OpenStudy (anonymous):

thanku!

ganeshie8 (ganeshie8):

np :) but read a bit on logifying... im sure u wil find somthing useful on that.. ppl wont simply put x = 1/(\lambda)^100.. and convert it to straight line.. . that doesnt help u in anyway....

ganeshie8 (ganeshie8):

i mean, u r not getting any benifit by simply putting x = 1/\lambda^2 or watever the graph looks straight line, but wat u wil be acheving by that ? you're masking the actual relatioonship of y = a + b/lambda^2

OpenStudy (anonymous):

lol yes practically not correct :P

ganeshie8 (ganeshie8):

is there any quantity in optics, that is related to wavelength by : q = k/lambda^2 if so, then i smell some use in putting simply x = 1/lambda^2 :)

OpenStudy (anonymous):

umm no :|

ganeshie8 (ganeshie8):

1/lambda = f right ?

OpenStudy (anonymous):

no :o

ganeshie8 (ganeshie8):

oops ! 1/T = f

OpenStudy (anonymous):

oh that lambda,yes

ganeshie8 (ganeshie8):

T = time period W = wavelength f = frequency

ganeshie8 (ganeshie8):

T : seconds W : meters f : hertz

ganeshie8 (ganeshie8):

W = lambda